iosangel:

A bicycle is classified by the diameter of its tires. For example, a 20-inch bicycle has tires that are 20 inches in diameter. The number of times a bicycle tire rotates in a given period of time is directly related to the distance traveled in that period of time. A 26-inch bicycle is pedaled so that the tires rotate at a rate of 200 revolutions per minute. A 20-inch bicycle is pedaled so that its tires rotate at a rate of 200 revolutions per minute, as well. What is the linear velocity of each bicycle? State your answers in inches per minute

Vocaloid:

for each bicycle, the wheel travels a linear distance equal to the circumference of the bicycle, per revolution. taking the 26-inch bicycle as an example, it will travel C = pi * d = pi * 26 inches per revolution. it travels 200 revolutions per minute, so you'd simply multiply the previous quantity by 200 to get the linear velocity in inches per minute. repeat this logic for the other bicycle.

iosangel:

Ok so for 2π x 20 x 26/2=1633.6*200=326720

Vocaloid:

Keep in mind that the 20 inch and 26 inch wheels belong to separate bicycles. For the 26 inch bicycle, the circumference would simply be 2pi * 26 / 2, there’s no 20 involved.

Luna13:

2π x 20 x 26/2=1633.6*200=326720

Vocaloid:

Like I *just* said, if you’re using 2pi r for the 26 inch wheel, it’s simply 2 pi * 26/2, you *don’t* multiply by 20 for the 26 inch wheel. The 20 inch belongs to a different wheel. The 20 would not be included in this calculation.

iosangel:

ok that wasnt my reply above but2pi * 6 /2=81.68 *200= 16336 then for 20 inch2 pi * 20/2= 62.83*200=12566

Wankle1222:

yes

Vocaloid:

Good